OpenStudy (anonymous):

Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. (5, 1), (–3, 4)

OpenStudy (karatechopper):

Ok. Well lets start by asking you this, do you know how to find the slope between two points?

OpenStudy (anonymous):

No not exactly..

OpenStudy (anonymous):

ok first off we need to use y2-y1/x2-x1 and its easy to figure out using it

OpenStudy (anonymous):

Ok. Thanks, I'll try to figure this out.

OpenStudy (anonymous):

and that formula is used to find out the slope

OpenStudy (anonymous):

This one will be a negative slope as well.

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) but we have to figure out the slope intercept form first for this one

OpenStudy (anonymous):

and then when you got both of these 2 then its easy to figure out standard form

OpenStudy (anonymous):

would the fraction be 2/5?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

umm no its going to be -3/8

OpenStudy (anonymous):

How did you solve that?

OpenStudy (anonymous):

using y2-y1/x2-x1

OpenStudy (anonymous):

Ok, I got it now.

OpenStudy (anonymous):

So now what?

OpenStudy (anonymous):

now we have the slope we figure out slope intercept form and you have to use graph paper for that

OpenStudy (anonymous):

plot both points and run a line through them

OpenStudy (anonymous):

and slope-intercept form is y=mx+b where m is the slope and b is the y intercept

OpenStudy (anonymous):

Can you help me out with that please?(:

OpenStudy (anonymous):

use graph paper and see what you get :)

OpenStudy (anonymous):

How do I graph it?

OpenStudy (anonymous):

ok take a peice of graph paper and make a large "+" in it and number the lines with 1's or 2's

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

but if number it with 1's you have to go like this 1,2,3,4,5 and so on and with 2's: 2,4,6,8 and so on

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

like this

OpenStudy (anonymous):

Ok I drew one.

OpenStudy (anonymous):

now did you plot both points on the graph?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

now go through the line until you see the line go through the "y" line

OpenStudy (anonymous):

x is the horizontal line on the "+" and y is vertical line of the "+"

OpenStudy (anonymous):

It passes through the Y line at 3.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now you got your slope intercept form and point slope form: y=-3/8x+3 and y-1=-3/8(x-5)

OpenStudy (anonymous):

now standard form is left

OpenStudy (anonymous):

well i got to go i hope you can figure this out